Owners of a recreation area are filling a small pond with water. They are adding water at a rate of
35
liters per minute. There are
700
liters in the pond to start.
Let W represent the amount of water in the pond (in liters), and let
T represent the number of minutes that water has been added. Write an equation relating
W to T, and then graph your equation using the axes below.
I am stuck on the graphing part. The graph shows the W = Y axis goes form 100-1,000 and the T = x axis goes from 1 -8.
How do I graph this?
Can you please help me.

Answer by ikleyn(53942)

.
Owners of a recreation area are filling a small pond with water. They are adding water at a rate of
35
liters per minute. There are
700
liters in the pond to start.
Let W represent the amount of water in the pond (in liters), and let
T represent the number of minutes that water has been added. Write an equation relating
W to T, and then graph your equation using the axes below.
I am stuck on the graphing part. The graph shows the W = Y axis goes form 100-1,000 and the T = x axis goes from 1 -8.
How do I graph this?
Can you please help me.
~~~~~~~~~~~~~~~~~~~~~~~~~

W = 700 + 35*T.



Plot y = 700 + 35*x